# A relativistic model of b-EMRI systems and their gravitational radiation

### v1

@@ Appendix D Conserved stress-energy constructed in the FFF
$z_{i}^{\mu}=\left(\tau,(-1)^{i}\frac{d}{2}\cos\omega_{IB}\tau,0,(-1)^{i}\frac{d}{2}\sin\omega_{IB}\tau\right).$ (91)
@@
$\frac{d\tau}{d\tau_{i}}=$ $\left(1-h_{\tilde{t}\tilde{t}}|_{z_{i}}-\left(\frac{d\omega_{IB}}{2}\right)^{2}\right)^{-2}+\mathcal{O}(d^{3}),$ (92)
having partially expanded in $d$ at fixed $\omega_{IB}$. Now consider the monopole stress-energy term
@@
for $\tilde{x}^{a}=(\tilde{x},\tilde{y},\tilde{z})$. Clearly this equation becomes trivially satisfied if the trajectories are instead geodesics of the Minkowski metric. We now expand the delta function in small $d$ at fixed $\omega_{IB}$;
@@
$P^{\mu}{}_{\nu}\ddot{z}_{i}^{\nu}\doteq-\omega_{IB}^{2}P^{\mu}{}_{\nu}z_{i}^{\nu},$ (100)
@@
is the required interaction term in the stress-energy to ensure its conservation though $\mathcal{O}(\mu d^{3})$ at fixed $\omega_{IB}$ having also introduced $z_{c}^{\alpha}\doteq\tau\tilde{t}^{\alpha}$ as the centre-of-mass worldline. Instead of using the projections, we could also write
@@
We can now compute the stress-energy of the IB in an arbitrary coordinate system. In hindsight, the form of the interaction term is not so mysterious — it is reminiscent of the Newtonian point particle rotational energy, $\mu\omega_{IB}^{2}r^{2}$, in which $r$ is the distance to the axis of rotation. As $d\rightarrow 0$, we recover the stress-energy of a free monopole particle of mass $\mu$. As a final remark, constructing such a conserved stress-energy tensor may seem arduous in the general case (un-equal mass eccentric IBs). However the relevant stress-energy tensor can also be constructed from the variation of the Lagrangian defined in Ref. [78, 77] (see e.g Chapter 4 of Ref. [115]).

### v2

@@ Appendix D Conserved stress-energy constructed in the FFF
$z_{i}^{\mu}=\left(\tau,(-1)^{i}\frac{d}{2}\cos\omega_{\mathrm{IB}}\tau,0,(-1)^{i}\frac{d}{2}\sin\omega_{\mathrm{IB}}\tau\right).$ (91)
@@
$\frac{d\tau}{d\tau_{i}}=$ $\left(1-h_{\tilde{t}\tilde{t}}|_{z_{i}}-\left(\frac{d\omega_{\mathrm{IB}}}{2}\right)^{2}\right)^{-2}+\mathcal{O}(d^{3}),$ (92)
having partially expanded in $d$ at fixed $\omega_{\mathrm{IB}}$. Now consider the monopole stress-energy term
@@
for $\tilde{x}^{a}=(\tilde{x},\tilde{y},\tilde{z})$. Clearly this equation becomes trivially satisfied if the trajectories are instead geodesics of the Minkowski metric. We now expand the delta function in small $d$ at fixed $\omega_{\mathrm{IB}}$;
@@
$P^{\mu}{}_{\nu}\ddot{z}_{i}^{\nu}\doteq-\omega_{\mathrm{IB}}^{2}P^{\mu}{}_{\nu}z_{i}^{\nu},$ (100)
@@
is the required interaction term in the stress-energy to ensure its conservation though $\mathcal{O}(\mu d^{3})$ at fixed $\omega_{\mathrm{IB}}$ having also introduced $z_{c}^{\alpha}\doteq\tau\tilde{t}^{\alpha}$ as the centre-of-mass worldline. Instead of using the projections, we could also write
@@
We can now compute the stress-energy of the IB in an arbitrary coordinate system. In hindsight, the form of the interaction term is not so mysterious — it is reminiscent of the Newtonian point particle rotational energy, $\mu\omega_{\mathrm{IB}}^{2}r^{2}$, in which $r$ is the distance to the axis of rotation. As $d\rightarrow 0$, we recover the stress-energy of a free monopole particle of mass $\mu$. As a final remark, constructing such a conserved stress-energy tensor may seem arduous in the general case (un-equal mass eccentric IBs). However the relevant stress-energy tensor can also be constructed from the variation of the Lagrangian defined in Ref. [80, 79] (see e.g Chapter 4 of Ref. [116]).